# Revisiting Einstein’s analogy: black holes as gradient-index lenses

### v1

@@ Abstract
According to Albert Einstein, gravitation is analogous to an optical medium. Building on this idea, various definitions of the gradient-index (GRIN) medium representing curved spacetime have been proposed; often, these approaches demand advanced knowledge of General Relativity and its associated mathematical methods. This paper introduces a novel approach for generating the form of GRIN media that reproduces the behavior of light in the presence of a Schwarzschild black hole or about the equatorial spacetime of a Kerr black hole. Our approach is based on Noether’s theorem that leads to optical Fermat’s principle, leveraging the system’s symmetry, resulting in a general method for finding GRIN media in a flat spacetime that reproduces the behavior of light in Schwarzschild-like spacetimes where only using photon trajectories as input. This approach is also valid for material particles and provides a complementary understanding of General Relativity.
@@ I Introduction
Gravitational lensing is the phenomenon where the trajectory of waves or particles is deflected due to the curvature of spacetime. Predicted by Einstein’s theory of General Relativity, this phenomenon has been explored through various spacetime solutions to Einstein’s field equations, demonstrating to be a powerful tool for studying the universe Grespan_2023 ; Rodr_guez_Fajardo_2023 .
@@
Many astrophysical studies rely on strong, weak, and microlensing phenomena to understand galaxy clusters’ structure and mass distribution. As natural cosmic lenses, these phenomena have led to the discovery of exoplanets beyond our solar system and have the potential to extend our observational capabilities beyond light. With advancements in gravitational wave detectors, the possibility of detecting lensed gravitational waves is increasing. For a comprehensive review of these subjects, refer to Perlick_2004 ; Rodr_guez_Fajardo_2023 ; Bartelmann_2010 ; Grespan_2023 .
@@
The deflection of light is often described using the thin-lens approximation. However, the direct detection of gravitational waves by the Laser Interferometer Gravitational-Wave Observatory (LIGO) Raab_2017 and the observation of the black hole at the center of our galaxy by the Event Horizon Telescope collaboration (EHT) Castelvecchi_2019 promote the use of the complete theory of General Relativity. Although phenomena involving the strong effects of curved spacetime are beyond our capacity to directly experience or experiment with, analogies with other branches of physics, such as classical mechanics, quantum mechanics, optics, and fluid mechanics Plebanski_1960 ; de_Felice_1971 ; Sch_tzhold_2002 ; Philbin_2008 ; Genov_2009 ; Chen_2010 ; leonhardt2010geometry ; Tinguely_2020 ; Rodr_guez_Fajardo_2023 ; _van_ara_2024 , may enable their study in the laboratory.
@@
Approaches to studying light propagation in curved spacetimes using concepts from optics and variational principles have been explored since the birth of general relativity. As noted in Ref. 8, ”the idea that gravitation is equivalent to an optical medium was first suggested by Einstein himself.” Key concepts along this path include Maxwell’s equations in general relativity as electrodynamics in a macroscopic medium Plebanski_1960 ; de_Felice_1971 , the optical metric Gordon_1923 , and a relativistic version of Fermat’s principle Kovner_1990 , among others.
@@
Metamaterials allow the reproduction of electromagnetic waves in curved spacetimes through the formal equivalence between Maxwell’s equations in general relativity and those of electrodynamics in a macroscopic medium. However, reproducing light rays through GRIN media is advantageous when polarization effects are disregarded, as it is much simpler and easier to construct experimentally Tinguely_2020 . Currently, various definitions of the gradient index for curved spacetime exist; therefore, as pointed out in Ref. 18, the best way to find the most precise definition is by comparing the null geodesics in spacetime with light rays in the optical media.
@@
Noether’s theorem, one of the most fascinating theorems in modern physics, was inspired by Albert Einstein’s work on relativity. It states that a natural law described by a Lagrangian function that exhibits symmetry has an invariant, also called conserved quantity. The associated Euler–Lagrange equations inherit the symmetry group of the variational problem kogan2003invariant . In general, variational principles are related to finding extrema, that in physical systems are commonly minima. In this sense, Fermat’s principle states that a ray of light traveling in a medium with variable refractive index will take the route that minimizes the time of travel between two given points, which can be expressed as a variational principle yielding to the corresponding Euler-Lagrange equations, simplifying substantially the problem.
@@
Suppose the Geometrical Optics formalism is used alone. In that case, finding the path of a ray of light in Gradient-Index (GRIN) media is cumbersome since it requires finding the solution of a partial differential equation called the Eikonal. Different methods have been proposed to solve this equation, but their implementation is complicated and is restricted to a medium with a specific GRIN distribution.
@@
In recent decades, the inverse problem—finding the GRIN distribution for a given light ray path—has garnered increasing attention. Since this task may be even more challenging than determining ray propagation in a GRIN medium, the methods proposed in the literature have different complexity. However, it has been demonstrated that an exact method for solving the inverse problem in any symmetric GRIN medium can be derived using Noether’s invariant of the system G_mez_Correa_2023 . This approach, known as the Physical GRIN Reconstruction (PhysGRIN) method, enables the determination of the refractive index at any point within the medium based on the ray direction and Noether’s invariant. Following the ray’s path and employing the invariant effectively defines the symmetric refractive index distribution. Once the ray trajectories are established, the algorithm is easily implemented and computationally efficient. This method can be extended to other areas of physics being formulated using variational principles, such as astronomy Schneider_1992 ; Price_2018 ; thorne2000gravitation ; landau1960mechanics ; wambsganss1998gravitational .
@@
In this paper, using the PhysGRIN method and leveraging the trajectory of a photon around the equatorial Kerr black hole, we introduce a novel approach for constructing GRIN lenses that replicate photon paths in the curved spacetime of these black holes. Inspired by Einstein’s analogy between gravitation and optical media, this method allows us to create GRIN models of curved spacetime without requiring advanced knowledge of general relativity or tensor calculus, instead relying on Noether’s theorem and the system’s symmetry.
II Photon trajectories around equatorial Kerr black holes
@@
The Schwarzschild metric is one of the most studied models of black holes; it is the unique, static, spherically symmetric vacuum solution to Einstein’s field equations and forms the basis for the theory of gravitational lensing Schneider_1992 . It is also a particular case of the Kerr spacetime, which generalizes it to include the effects of rotation.
@@
The infinitesimal interval of events for a Kerr black hole of mass $M$ in Boyer–Lindquist coordinates is given by
$ds^{2}=-c^{2}\left(1-\frac{2GMr_{bl}}{c^{2}\Sigma}\right)dt^{2}-\frac{4GMr_{bl}a\sin^{2}\theta}{c^{2}\Sigma}cdtd\phi$
@@
$+\left(r_{bl}^{2}+a^{2}+\frac{2GMr_{bl}a^{2}}{c^{2}\Sigma}\sin^{2}\theta\right)\sin^{2}\theta d\phi^{2}+\frac{\Sigma}{\Delta}dr_{bl}^{2}+\Sigma d\theta^{2},$ (1)
@@
where $G$ denotes the universal gravitational constant, $c$ the speed of light, $a$ is related to the angular momentum $J$ of the black hole by $a=J/Mc$, $(r_{bl},\theta,\phi)$ are related to the Cartesian coordinates by $x_{c}=\sqrt{r_{bl}^{2}+a^{2}}\sin\theta\cos\phi$, $y_{c}=\sqrt{r_{bl}^{2}+a^{2}}\sin\theta\sin\phi$, and $z_{c}=r_{bl}\cos\theta$, and
@@
This metric has two coordinate singularities determined by $\Delta=0$. The exterior singularity, $r_{bl,+}=\left[GM+\sqrt{G^{2}M^{2}-c^{4}a^{2}}\right]/c^{2}$, is known as the outer event horizon and marks the boundary beyond where nothing can escape. The interior singularity, $r_{bl,-}=\left[GM-\sqrt{G^{2}M^{2}-c^{4}a^{2}}\right]/c^{2}$, is called the inner event horizon or Cauchy horizon, and marks the region where spacetime becomes unstable mcnamara1978instability . These coordinate singularities restrict the black hole’s angular momentum, which must satisfy $-GM/c\leq a\leq GM/c$ for $r_{bl,+}$ and $r_{bl,-}$ to be real.
The Kerr metric also contains a true singularity at $r_{bl}=0$ and $\theta=\pi/2$, where $\Sigma=0$, causing the curvature to become infinite. This singularity forms a ring of radius $a$ at the equator. Unlike the Schwarzschild metric $(a=0)$, the coefficient of $dt^{2}$ in the Kerr metric changes sign at a region different from the event horizon, known as the ergosphere. The boundary of the ergosphere is given by $r_{bl,e}=\left[GM+\sqrt{G^{2}M^{2}-a^{2}c^{4}\cos^{2}\theta}\right]/c^{2}$, which is exterior to the event horizon except at the poles $(\theta=0,\pi)$, where they coincide. Inside the ergosphere, static particles $(dr_{bl}=d\theta=d\phi=0)$ are not possible because their worldlines would be forced to move in the direction of the black hole’s angular momentum due to the dragging effect of spacetime.
@@
$ds^{2}=g_{\alpha\beta}dx^{\alpha}dx^{\beta},$ (3)
where the upper and lower indices denote Einstein’s summation notation, with $\alpha,\beta=0,1,2,3$, and $x^{0}=ct$, $x^{1}=r$, $x^{2}=\theta$, and $x^{3}=\phi$. The world line of a particle, which represents the path that the particle traces in spacetime, is a one-dimensional curve in four-dimensional spacetime. This curve can be parametrized by a parameter $\lambda$, which labels every point along the trajectory. For massive particles, it is convenient to identify $\lambda$ with the particle’s proper time $\tau$. However, proper time is useless for massless particles like photons, so $\lambda$ is used instead. The parameterization of the world line is denoted as
@@
The standard approach to null geodesics is through the Lagrangian formalism, though they can also be expressed in the Hamiltonian formalism. A brief discussion on these formalisms and the challenges of computing particle or photon trajectories can be found in Ref. 22. Since we are interested in geodesics confined to the equatorial plane to exploit spherical symmetry for GRIN medium reconstruction, the most straightforward approach is via Lagrangian formalism. The Lagrangian associated with the metric (3) is given by:
@@
where $\dot{x}^{\alpha}=dx^{\alpha}/d\lambda$ denotes the components of the four-velocity. The term $g(\dot{x},\dot{x})\equiv g_{\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\beta}$ represents the square of the norm of the four-velocity, which for photons equals zero, as the four-velocity vector lies on the local light cone. Under these conditions, null geodesics (world lines of photons) satisfy the following system of equations:
$\mathcal{L}=0,\quad\frac{d}{d\lambda}\frac{\partial\mathcal{L}}{\partial\dot{x}^{\alpha}}-\frac{\partial\mathcal{L}}{\partial x^{\alpha}}=0.$ (6)
A review and classification of trajectories using elliptic integrals of the first and third kind, Carlson’s elliptic integrals, Jacobi-elliptic functions, and Weierstrass elliptic functions, along with codes for computing geodesics in Wolfram Mathematica and Fortran, can be found in Ref. 27; 28; 29; 30; 31; 32.
For convenience, let us define $m_{b}\equiv GM/c^{2}$ and $T=ct$. The geodesics lying in the equatorial plane must satisfy $\theta=\pi/2$ and $\dot{\theta}=0$. A direct insight is that the Kerr metric (and thus the Lagrangian) does not depend explicitly on $t$ and $\phi$. Therefore, we have the following two integrals Chandrasekhar_1998 :
@@
where $E$ and $L$ are the energy per unit mass and the angular momentum per unit mass, respectively. Thus, $\dot{T}$ and $\dot{\phi}$ can be determined in terms of $r_{bl}$, $L$, and $\varepsilon$, and substituted into $ds^{2}=0$ (the condition for null geodesics, equivalent to $\mathcal{L}=0$), and an equation for $\dot{r}_{bl}$ is derived Chandrasekhar_1998 .
@@
Since we are only interested in the trajectories of photons, we can eliminate the dependence on the affine parameter as $\dot{\phi}/\dot{r}_{bl}=d\phi/dr_{bl}$, which can be rewritten as fan2018analytical
$d\phi=\frac{hr_{bl}^{2}-2m_{b}r_{bl}(h-a)}{\Delta\sqrt{N}}\,dr_{bl},$ (9)
@@
When $h=a$, the integration of Eq. (9) is straightforward Chandrasekhar_1998

### v2

@@ Abstract
According to Albert Einstein, gravitation is analogous to an optical medium. Building on this idea, various definitions of the gradient-index (GRIN) medium representing curved spacetime have been proposed; often, these approaches demand advanced knowledge of General Relativity (GR) and its associated mathematical methods. This paper introduces an alternative approach for generating the form of GRIN media that reproduces the exact behavior of photon trajectories in the presence of a Schwarzschild black hole or about the equatorial plane of a Kerr black hole. Our approach is based on Noether’s theorem that leads to the optical Fermat principle, leveraging the system’s symmetry, resulting in an unrestricted and robust method for finding GRIN media in a flat spacetime that reproduces the behavior of null geodesics in Schwarzschild-like spacetimes, employing only photon trajectories as input. This approach is also valid for massive particles and provides a complementary understanding of GR. For this reason, the method is applied to generate the GRIN medium from the S2 star’s trajectory around SgrA*, demonstrating that sufficient precision can be achieved to reproduce the orbit’s subtle apsidal precession.
@@ I Introduction
Gravitational lensing is the phenomenon in which waves (including gravitational waves) or particles are deflected due to the curvature of spacetime. The idea that gravity affects light has been considered for a long time. In the first edition of Opticks (1704), Isaac Newton suggested that bodies could influence the behavior of light, implying a deflection of its rays Schneider_1992 . In 1801, Johann Georg von Soldner published a quantitative analysis of this effect within the framework of Newtonian mechanics. Independently, Albert Einstein first described the effect of gravitation on light in 1911 within the context of special relativity Einstein_1911 . Four years later, he extended his analysis to the general theory of relativity (GR), thus incorporating the curvature of spacetime and providing a precise quantitative prediction of light deflection einstein1915perihelion . Since then, gravitational lensing has been extensively studied through various spacetime solutions to Einstein’s field equations, proving to be a powerful tool for exploring the universe.
@@
Many astrophysical studies rely on strong, weak, and microlensing phenomena to understand galaxy clusters’ structure and mass distributions. As natural cosmic lenses, these phenomena have led to the discovery of exoplanets beyond our solar system and have the potential to extend our observational capabilities beyond light. With advancements in gravitational wave detectors, the possibility of detecting lensed gravitational waves is increasing. For a comprehensive review of these subjects, we refer the interested reader to Perlick_2004 ; Rodr_guez_Fajardo_2023 ; Bartelmann_2010 ; Grespan_2023 .
The deflection of light is often described using the thin-lens approximation. However, the direct detection of gravitational waves by the Laser Interferometer Gravitational-Wave Observatory (LIGO) Raab_2017 and the observation of the black hole shadow at the center of our galaxy by the Event Horizon Telescope collaboration (EHT) Castelvecchi_2019 promote the use of the complete theory of GR. Efforts in this direction include the Virbhadra-Ellis lens equation model Virbhadra_2000 , which is valid for large deflection angles of rays, and the description of other effects Virbhadra_2009 ; Virbhadra_2022 ; Virbhadra_2024 ; Virbhadra_2024_c .
Although the black hole shadow is already a manifestation of strong gravitational fields, next-generation telescopes such as the ngEHT Johnson_2023 will improve our ability to observe additional relativistic features that remain unresolved with current instrumentation. These include the photon ring, the inner shadow, and dynamic signatures such as light echoes and polarization structures near the event horizon. For a historical overview of gravitational lensing, the reader is referred to the foundational works already mentioned above. For comprehensive reviews of gravitational lensing theory and its applications in astrophysics and cosmology, see Perlick_2004 ; Bartelmann_2010 . For more recent studies on strong lensing by Schwarzschild black holes, including image distortions and relativistic images, see Virbhadra_2000 ; Virbhadra_2009 ; Virbhadra_2022 ; Virbhadra_2024 ; Virbhadra_2024_c .
@@
Although phenomena involving the strong effects of curved spacetime are beyond our capacity to directly experience or experiment with, analogies with other branches of physics, such as classical mechanics, quantum mechanics, optics, and fluid mechanics Plebanski_1960 ; de_Felice_1971 ; Sch_tzhold_2002 ; Philbin_2008 ; Genov_2009 ; Chen_2010 ; leonhardt2010geometry ; Tinguely_2020 ; Rodr_guez_Fajardo_2023 ; _van_ara_2024 , may enable their analogue simulations in the laboratory.
@@
Approaches to studying light propagation in curved spacetimes using concepts from optics and variational principles have been explored since the birth of GR eddington1920report ; eddington1921space . As noted in Ref. de_Felice_1971 , “the idea that gravitation is equivalent to an optical medium was first suggested by Einstein himself.” Key concepts along this path include Maxwell’s equations in GR as electrodynamics in a macroscopic medium, the optical metric, and a relativistic version of Fermat’s principle , among others Plebanski_1960 ; de_Felice_1971 ; Gordon_1923 ; Kovner_1990 .
@@
Metamaterials allow the reproduction of the behavior of the electromagnetic waves in curved spacetimes through the formal equivalence between Maxwell’s equations in GR and those of electrodynamics in a macroscopic medium. However, reproducing light rays through GRIN media is advantageous when polarization effects are neglected, as it is much simpler and easier to construct experimentally Tinguely_2020 . Currently, various definitions of the gradient index for curved spacetimes that mimics photon trajectories exist; therefore, following the comments in Refs. Parvizi_2024 ; Ramezani_Aval_2024 , the core of the matter is how to design optical devices or metamaterials that exactly reproduce photon trajectories.
@@
Noether’s theorem, one of the most fascinating theorems in modern physics, was inspired by Albert Einstein’s work on relativity. It states that a natural law described by a Lagrangian function that exhibits symmetry has an invariant, a conserved quantity. The associated Euler–Lagrange equations inherit the symmetry group of the variational problem kogan2003invariant . In general, variational principles are related to finding extrema that, in physical systems, are commonly minima. In this sense, Fermat’s principle states that a ray of light traveling in a medium with variable refractive index will take the route that minimizes the time of travel between two given points, which can be expressed as a variational principle yielding to the corresponding Euler-Lagrange equations, substantially simplifying the problem.
@@
Suppose the geometrical optics formalism is used alone. In that case, finding the path of a ray of light in Gradient-Index (GRIN) media is cumbersome since it requires finding the solution of a partial differential equation called the Eikonal. Different methods have been proposed to solve this equation, but their implementation is complicated and is restricted to a medium with a given GRIN distribution.
@@
In recent decades, the inverse problem—finding the GRIN distribution for a given light path—has garnered increasing attention. Since this task may be even more challenging than determining ray propagation in a GRIN medium, the methods proposed in the literature have different complexities. However, it has been demonstrated that an exact method for solving the inverse problem in any symmetric GRIN medium can be derived using Noether’s invariant of the system G_mez_Correa_2023 . This approach, known as the Physical GRIN Reconstruction (PhysGRIN) method, enables the determination of the refractive index at any point within the medium based on the ray direction and Noether’s invariant. Following the ray’s path and employing the invariant effectively defines the symmetric refractive index distribution. Once the ray trajectories are established, the algorithm is easily implemented and computationally efficient. This method, being formulated using the variational principle, can be extended to other areas of physics, such as astronomy Schneider_1992 ; Price_2018 ; thorne2000gravitation ; landau1960mechanics ; wambsganss1998gravitational .
@@
In this paper, using the PhysGRIN method and leveraging the trajectory of a photon in the equatorial plane of a Schwarzschild black hole or in the equatorial plane of a Kerr black hole, we introduce an alternative approach for constructing GRIN lenses that replicate photon paths in the curved spacetime of these black holes. Inspired by Einstein’s analogy between gravitation and optical media, this method allows us to create GRIN models of curved spacetimes without requiring advanced knowledge of GR or tensor calculus, instead relying on Noether’s theorem and the system’s symmetry. The robustness of the PhysGRIN method is tested by adding noise to exact null geodesics, followed by a discussion on how to best handle stochastic data. Since PhysGRIN relies solely on a given trajectory, it can also be applied to massive particles. As a case study, the gradient index profile for the orbit of the S2 star around SgrA*, the nearest candidate supermassive black hole, is reconstructed, demonstrating that the method is both simple and accurate enough to reproduce its subtle apsidal precession 2020 .
II Photon trajectories in the equatorial plane of Kerr black holes
@@
The Schwarzschild metric is one of the most studied models of black holes; it is the unique, static, spherically symmetric vacuum solution to Einstein’s field equations and provides a fundamental framework for investigating gravitational lensing (see Schneider_1992 ; Virbhadra_2009 , and references therein). It is also a particular case of the Kerr spacetime, which generalizes it to include the effects of rotation.
@@
The line element for a Kerr black hole of mass $M$ in Boyer–Lindquist coordinates is given by
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}s^{2}=-c^{2}\left(1-\frac{2GMr_{bl}}{c^{2}\Sigma}\right){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}t^{2}-\frac{4GMr_{bl}a\sin^{2}\theta}{c^{2}\Sigma}c{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}t{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}\phi$
@@
$+\left(r_{bl}^{2}+a^{2}+\frac{2GMr_{bl}a^{2}}{c^{2}\Sigma}\sin^{2}\theta\right)\sin^{2}\theta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}\phi^{2}+\frac{\Sigma}{\Delta}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}r_{bl}^{2}+\Sigma{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}\theta^{2},$ (1)
@@
where $G$ denotes the universal gravitational constant, $c$ the speed of light, $a$ is related to the angular momentum $J$ of the black hole by $a=J/Mc$, and
@@
This metric has two coordinate singularities determined by $\Delta=0$. The exterior singularity, $r_{bl,+}=\left[GM+\sqrt{G^{2}M^{2}-c^{4}a^{2}}\right]/c^{2}$, is known as the outer event horizon and marks the boundary beyond where nothing can escape. The interior singularity, $r_{bl,-}=\left[GM-\sqrt{G^{2}M^{2}-c^{4}a^{2}}\right]/c^{2}$, is called the inner event horizon or Cauchy horizon, and marks the region where the spacetime becomes unstable mcnamara1978instability . These coordinate singularities restrict the black hole’s angular momentum, which must satisfy $-GM/c\leq a\leq GM/c$ for $r_{bl,+}$ and $r_{bl,-}$ to be real.
The Kerr metric also contains a true singularity at $r_{bl}=0$ and $\theta=\pi/2$, where $\Sigma=0$, causing the curvature to become infinite. It is a common depiction to represent such a singularity as a ring of radius $a$ in the equatorial plane. This follows from the relationship between Boyer-Lindquist and Cartesian coordinates: $x_{c}=\sqrt{r_{bl}^{2}+a^{2}}\sin\theta\cos\phi$, $y_{c}=\sqrt{r_{bl}^{2}+a^{2}}\sin\theta\sin\phi$, and $z_{c}=r_{bl}\cos\theta$. For this reason, from now on, we will visualize the null geodesics graphically in Cartesian coordinates.
Unlike the Schwarzschild metric $(a=0)$, the coefficient of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}t^{2}$ in the Kerr metric changes sign at a region different from the event horizon, known as the ergoregion. The boundary of the ergoregion (ergosurface) is given by $r_{bl,e}=\left[GM+\sqrt{G^{2}M^{2}-a^{2}c^{4}\cos^{2}\theta}\right]/c^{2}$, which is exterior to the event horizon except at the poles $(\theta=0,\pi)$, where they coincide. Inside the ergoregion, static particles $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}r_{bl}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}\theta={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}\phi=0)$ cannot exist because their worldlines would be forced to move in the direction of the black hole’s angular momentum due to the frame-dragging effect.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}s^{2}=g_{\alpha\beta}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}x^{\alpha}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}x^{\beta},$ (3)
where the upper and lower indices denote Einstein’s summation notation, with $\alpha,\beta=0,1,2,3$, and $x^{0}=ct$, $x^{1}=r$, $x^{2}=\theta$, and $x^{3}=\phi$. The world line of a particle, which represents the path that the particle traces in spacetime, is a one-dimensional curve in four-dimensional spacetime. This curve can be parameterized by a parameter $\lambda$, which labels each point along the trajectory. For massive particles, it is convenient to identify $\lambda$ with the particle’s proper time $\tau$. However, since proper time is not defined for massless particles such as photons, $\lambda$ is instead interpreted as an affine parameter. The parameterization of the world line is denoted as
@@
The standard approach to null geodesics is through the Lagrangian formalism, though they can also be expressed in the Hamiltonian formalism. A brief discussion on these formalisms and the challenges of computing particle or photon trajectories can be found in Ref. Price_2018 . Since we are interested in geodesics confined to the equatorial plane to exploit spherical symmetry for GRIN medium reconstruction, the most straightforward approach is via the Lagrangian formalism. The Lagrangian associated with the metric (3) is given by:
@@
where $\dot{x}^{\alpha}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}x^{\alpha}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}\lambda$ denotes the components of the four-velocity. The term $g(\dot{x},\dot{x})\equiv g_{\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\beta}$ represents the square of the norm of the four-velocity, which for photons equals zero, as the four-velocity vector lies on the local light cone. Under these conditions, null geodesics (world lines of photons) satisfy the following system of equations:
$\mathcal{L}=0,\quad\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}\lambda}\frac{\partial\mathcal{L}}{\partial\dot{x}^{\alpha}}-\frac{\partial\mathcal{L}}{\partial x^{\alpha}}=0.$ (6)
A review and classification of trajectories using elliptic integrals of the first and third kind, Carlson’s elliptic integrals, Jacobi-elliptic functions, and Weierstrass elliptic functions, along with codes for computing geodesics in Wolfram Mathematica and Fortran, can be found in Ref. Dexter_2009 ; Yang_2013 ; fan2018analytical ; Gralla_2020 ; Cie_lik_2023 ; Liu__2024 .
At this point, it is important to clarify that this work aims to demonstrate the simplicity and accuracy of PhysGRIN in reconstructing a GRIN medium that reproduces photon trajectories, which can be exploited without major complications when light trajectories exhibit symmetries. The method used to calculate the null geodesics (whether analytical, numerical, or approximate) is independent of PhysGRIN. Therefore, in what follows, we present specific analytical results in the simplest possible form to validate the subsequent numerical integration of photon trajectories, as this approach does not require more elaborate mathematical notions or representations, which are beyond the scope of the present work.
For convenience, let us define $m_{b}\equiv GM/c^{2}$ and $T=ct$. The geodesics lying in the equatorial plane must satisfy $\theta=\pi/2$, $\dot{\theta}=0$, and $\ddot{\theta}=0$. A direct insight is that the Kerr metric (and thus the Lagrangian) does not depend explicitly on $t$ and $\phi$. Therefore, we have the following two integrals Chandrasekhar1983mathematical :
@@
where $E$ and $L$ denote the energy per unit mass and the angular momentum per unit mass, respectively. Thus, from Eqs. (7), $\dot{T}$ and $\dot{\phi}$ can be expressed in terms of $r_{bl}$, $L$, and $\varepsilon$. Substituting these relations into the null condition ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}s^{2}=0$ (equivalently, $\mathcal{L}=0$), one obtains the following equation for $\dot{r}_{bl}$ Chandrasekhar1983mathematical :
@@
Since we are only interested in the trajectories of photons, we can eliminate the dependence on the affine parameter as $\dot{\phi}/\dot{r}_{bl}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}\phi/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}r_{bl}$, which can be rewritten as fan2018analytical
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}\phi=\frac{hr_{bl}^{2}-2m_{b}r_{bl}(h-a)}{\Delta\sqrt{N}}\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}r_{bl},$ (9)
@@
When $h=a$, the integration of Eq. (9) is straightforward Chandrasekhar1983mathematical